OpenStudy (gina):

find the general solution of the equation,,,y'+2y=cosx

OpenStudy (anonymous):

Oh dang...i got to brush up on differential equations...um...what did you guys recently learn? What methods?

OpenStudy (anonymous):

it's the same as the other one i told you, linear first order

OpenStudy (gina):

just show me your answer just wanna be sure

OpenStudy (anonymous):

well, you can always differentiate and plug it back into the original to check if you want with something simple like this, but anyhow, http://www.wolframalpha.com/input/?i=y%27%2B2y%3Dcosx wolfram will make life so much easier

OpenStudy (gina):

\[(x primey -y)arctgy divx =x; y \x=1=0\]

OpenStudy (gina):

\[spaceknight i got a diferent answre and my answer \in the book is ,,,,x^2 +y^2=lncx ^{2}\]

OpenStudy (anonymous):

wolfram doesn't lie so either your book is wrong or you're looking at the wrong answer

OpenStudy (gina):

\[y'=10^{x+y}\]

OpenStudy (gina):

thanx can i also solve physics on that site?

OpenStudy (anonymous):

Gina, did you manage to solve this..... also what is the answer in your book?

OpenStudy (gina):

my answer is x^2+y^2=\[\ln cx ^{x}\]

OpenStudy (gina):

but did not manage it(

OpenStudy (anonymous):

This is my answer\[y=2/5*\cos(x)+1/5*\sin(x)+c\]

OpenStudy (gina):

\[but how from this ,,,y'=10^{x+y}??\]

OpenStudy (anonymous):

nadeem you forgot to divide the e^2x into the constant, but rest is right.

OpenStudy (anonymous):

sorry typo this would be it: \[y=2/5∗\cos(x)+1/5∗\sin(x)+ce^{2x}\]

OpenStudy (anonymous):

Thanks for the reminder spaceknight

OpenStudy (anonymous):

Solve the differential equation using an integrating factor, which is 2 in your case

OpenStudy (anonymous):

not 2 but e^2x

OpenStudy (gina):

help with physics ,,,,the space between the plates of a plane capacitor filled with mica (e = 7). area of the capacitor plates is 50cm ^ 2. determine the surface density of bound charges on mica, if the capacitor plates attract each other with a force

OpenStudy (gina):

\[\int\limits_{?}^{?} dx \div cosx \times \sin ^{3}x\]